Use the data provided to find the proportion of respondents who own​ Dell, Apple, and HP computers.

Brand Proportion
Dell ?
Apple ?
HP ?
Brand Cost Year of Purchase Purchase reason Rating (1 low, 5 high)
Dell 700 2014 Cost 4
Apple 1,500 2014 Brand name 4
Dell 800 2014 Cost 5
Dell 800 2012 Cost 4
Dell 1,000 2011 Cost 4
Apple 1,000 2012 Brand name 5
Apple 1,200 2012 Prior use 4
Dell 1,000 2010 Cost 3
Dell 800 2012 Brand name 5
Dell 700 2011 Customer service 5
Apple 1,300 2012 Brand name 4
HP 600 2013 customer service 5
Answer #1

From the table it can be seen that:

Total No. of Respondents = 12

No. of respondents who own​ Dell = 7

So, proportion of respondents who own​ Dell = 7/12 = 0.5833

No. of respondents who own​ Apple = 4

So, proportion of respondents who own​ Apple = 4/12 = 0.3333

No. of respondents who own​ HP = 1

So, proportion of respondents who own​ HP = 1/12 = 0.0833

Brand Proportion
Dell 0.5833
Apple 0.3333
HP 0.0833